# : Context Management, Validation, and Transaction Guarantees for Multi-Agent LLM Planning

### v1

@@ 3.2. Transaction State Management
addresses the challenge of enforcing ACID guarantees in a distributed LLM agent environment with stateless agents. implements spatial-temporal state identification through a comprehensive transaction state management approach that preserves selective context information throughout multi-agent workflows. Using the MACI algorithm (Chang, 2024a) that we have previously developed, the states and dependencies of a given application specification can be automatically extracted to enable effective transaction state management.
@@ 3.2.1. State Dimensions and Context Preservation
maintains three critical state dimensions to support both transaction management and context preservation:
@@
Application State (): The application-domain dependent data being manipulated (e.g., travel booking details).
@@
Operation State (): Records of executed operations, including inputs, outputs, and effects on domain entities.
@@
Dependency State (): An explicit mapping of inter-operation dependencies (constraints) and their satisfaction criteria.
By structuring operations, dependencies, and application data, ensures that essential information remains accessible even as LLMsβ attention mechanisms might otherwise cause older context to be lost. TableΒ 1 details the specific information recorded and mechanisms used to implement these state dimensions. Note that the decision reasoning recorded under is a critical component where LLMs make significant contributions that traditional workflow engines cannot support.
For example, in the Thanksgiving dinner planning scenario from SectionΒ 5.2, would maintain application state for each family memberβs location and activities, operation state for travel arrangements and food preparation steps, and dependency state capturing constraints like βturkey must be in the oven by 2:00 PMβ and βJames must pick up Emily after exiting the terminal.β
@@ 3.2.2. Dependency Tracking for Compensation
maintains formal representations of operation dependencies in a directed graph structure:
where represents the condition that must be satisfied for the dependency to be resolved. These conditions can be complex boolean expressions combining multiple upstream requirements, such as:
where represents an arbitrary Boolean function that determines how the upstream conditions () are combined.
This dependency graph is crucial for compensation planning during failures. When an operation fails, traverses the dependency graph to identify the operations that may be affected, creating a precise compensation sequence that ensures that the consistency of the system is maintained. The graph also enables to identify and preserve only the critical dependencies needed for potential compensation, reducing overhead.
@@ 3.2.3. Intra-Agent Critical Context Tracking
preserves critical context that LLMs might otherwise forget during extended operations. This includes agent specifications, intermediate reasoning states (chain-of-thought), decision justifications, and the context required for potential compensation actions. Empirical study (Xiao etΒ al., 2024; Modarressi etΒ al., 2025; Liu etΒ al., 2024a; Hsieh etΒ al., 2024) shows that LLMs can forget simple facts explicitly provided in context, such as travel times between locations.
The system implements a validation cycle where agent operations produce selected logged context that undergoes validation against system specifications. Valid operations proceed to the next workflow node, while invalid ones trigger compensating actions before retrying. This ensures that:
@@
Compensation mechanisms have necessary historical context.
@@
Operations remain consistent with defined specifications.
@@
Dependency relationships maintain integrity throughout the workflow.
This context tracking directly addresses the context narrowing issue observed in our experiments in SectionΒ 5.2, where models like Claude 3.7 lost track of critical constraints such as fire safety and travel times during reactive planning.
@@ 3.2.4. Inter-Agent Communication Protocol
implements a structured communication protocol between agents that explicitly manages what information is passed between them. This protocol ensures that:
@@
Each transaction agent receives precisely the information needed for its operation.
@@
Dependencies and constraints are properly communicated.
@@
boundaries are respected in multi-agent operations.
The protocol explicitly separates the agentβs working context (input prompt, reasoning space) from the transaction state. This separation allows agents to operate with minimal context while the system maintains complete transaction information, addressing scalability concerns in complex multi-agent workflows.
@@ 3.3. Independent Validation Framework
With transaction state management ensuring persistence, we now examine how implements its independent validation of critical junctures to address the LLM self-validation limitations described in SectionΒ 1.
@@ 3.3.1. Multi-Level Validation Architecture
implements a two-tier validation approach that separates intra-agent and inter-agent concerns.
@@
This validation operates within an individual LLM agent, addressing both theoretical limitations (GΓΆdelβs incompleteness) and practical challenges (attention narrowing, context forgetting).
Syntactic Validation: Ensures outputs conform to expected formats, structures, and schemas (e.g., JSON validity).
Semantic Validation: Verifies logical consistency, coherence, and significance of the agentβs operations.
Factual Validation: Confirms that critical facts provided in context are maintained throughout the reasoning process (e.g., travel times, constraints).
Constraint Adherence: Verifies that agent decisions respect all explicitly defined constraints from specifications.
Reasoning Validation: Evaluates whether the agentβs chain-of-thought process follows logically sound inference patterns.
Context Retention: Confirms the agent has maintained necessary contextual information required for its current operation.
In the Thanksgiving dinner scenario, this validation would catch the fire safety violation where Sarah was scheduled to leave while the turkey was unattended in the oven.
@@
This validation ensures proper information flow and dependency management across multiple agents in a workflow.
Contract Validation: Verifies that inputs and outputs between agents conform to predefined interfaces.
Dependency Satisfaction: Confirms that all prerequisites are met before an agent operation is executed.
Consistency Checks: Ensures shared data elements maintain consistent meanings across agent boundaries.
Temporal Validation: Verifies that operations happen in the correct sequence, respecting causal dependencies.
Mutual Agreement: For collaborative decisions, confirms that multiple agents have reached consensus on shared facts.
Transaction Boundary Integrity: Ensures that transaction ACID properties are maintained across agent operations.
For example, in the wedding coordination scenario, inter-agent validation would ensure that Patβs position is consistently tracked across agents, preventing the error where DeepSeek R1 βrewrote historyβ by starting Patβs drive to the airport after the traffic alert despite Pat already being at the airport.
@@ 3.3.2. Validation Protocols and Response Mechanisms
To operationalize the validation types in TableΒ 2, implements structured protocols that govern system responses to validation outcomes:
Rejection Protocol: Triggers compensating transactions upon validation failure (e.g., canceling hotel bookings if flight validation fails).
Augmentation Protocol: Refines agent outputs while preserving agent autonomy (e.g., formatting restaurant recommendations without changing selections).
Feedback Protocol: Captures validation results to improve future agent operations (e.g., noting that travel time estimates were consistently underestimated).
@@ 3.4. Failure Handling and Recovery
βs failure handling leverages the transaction state and dependency tracking described in the previous sections to implement comprehensive recovery:
Operation-Level Recovery: When individual operations fail, compensating actions execute to semantically undo their effects.
Workflow-Level Recovery: When entire workflows fail, the system orchestrates multiple compensations in the correct order based on the captured dependency graph.
For each compensating transaction that reverses a failed transaction , the system leverages the preserved transaction state to provide the original inputs and context that received, the outputs and side effects produced, and the specific aspects of that must be undone. This approach ensures reliable compensation even when the original agent is no longer available or when the operation occurred many steps earlier in a complex workflow.
@@ 4. Design with Travel Planning
This application helps users plan international trips with multiple destinations and manage complex itineraries within a specified budget. The system leverages LLMs to generate and refine feasible travel plans, followed by a transactional booking process implemented with to ensure consistency.
@@ 4.2. System Workflow
The application workflow is divided into two main phases: itinerary planning and booking execution. FigureΒ 2 depicts the work flow with phase one in gray and phase two in green color.
@@ 4.2.1. Phase 1: Itinerary Planning
LLM generates multiple feasible itineraries based on user requirements
@@
Each itinerary includes flight options, hotel suggestions, and transportation (train) between cities
@@
System estimates costs for each component and validates against budget constraints
@@
User reviews suggested itineraries and provides feedback
@@
LLM adjusts plans based on feedback while maintaining budget constraints
@@
System tracks and preserves critical context (travel dates, preferences, constraints)
@@
Process repeats until user is satisfied
@@
User selects a final itinerary
@@
System generates a detailed breakdown of all necessary bookings
@@
Critical dependencies between bookings are identified (e.g., hotel check-in must be after flight arrival)
@@ 4.2.2. Phase 2: Booking Process ( Implementation)
Booking Workflow Creation:
@@
System constructs a SAGA transaction sequence based on the finalized itinerary
@@
Defines compensation actions for each booking step
@@
Establishes validation rules for each transaction
@@
International Flight Booking: SFO Berlin
@@
Berlin Hotel Booking
@@
Train Booking: Berlin Cologne
@@
Cologne Hotel Booking
@@
International Flight Booking: Cologne SFO
@@
Potential failures:
Flight unavailability or price changes
Hotel rooms no longer available
Train scheduling conflicts
@@
Compensation strategies:
Cancel previous bookings (with cancellation policies)
Replan affected portions of the itinerary
Suggest alternatives within similar price range
@@ 5.1. Experimental Design
We evaluated four LLMsβClaudeΒ 3.7, DeepSeekΒ R1, GPT-4o, and GPT-o1βalongside our proposed framework. All experiments were conducted between MarchΒ 12 andΒ 17, 2025.
@@ 5.2.2. Context Narrowing
When tested with GPT-o1, all constraints were correctly observed. However, at the final step, it added 30 extra minutes to Jamesβ driving time from Boston Airport to home, citing potential traffic congestion. This kind of βclevernessβ is concerning, as an LLM may inject its own opinions at unpredictable stages in unpredictable ways. In this case, we had to explicitly prompt GPT-o1 to disclose its βcommon senseβ assumptions for verification.
@@
Remediation: Context Management
and Reactive Planning
To remedy the problem of context narrowing/loss, employs a context management agent to checkpoint historical state transitions, unresolved dependencies, and constraints. A key design criteria is to keep the agentβs context small, to avoid its own sifferring from attention narrowing. Hence, we employ two validation agents, one for travel coordination and the other for food preparation, each keeps a context of less than .
For problem P9, we instruct the agent to store each individualβs current state, next scheduled state-transition time, and all relevant scheduling constraints.
When an unexpected event triggers reactive planning, all individuals roll back to the last saved state. The system then consolidates past and new constraints, resolving conflicts before proceeding with planning. This ensures that:
Past history is not inadvertently overridden.
New dependencies and constraints are properly integrated.
The system maintains consistency across state transitions.
@@ 5.3.2. Transaction Properties
Alert 1:00 PM: Traffic Alert, an accident near the Logan airport in Boston triples all travel times to and from the airport!
@@
Below is a list of results from four LLMs:
@@
DeepSeek R1: TableΒ 6 demonstrates how DeepSeekΒ R1 fails to maintain temporal consistency in reactive planning. When the traffic alert takes effect at 1:00βPM, DeepSeek discards its execution history and attempts to create a new plan starting from that point onward. Critically, it reassigns Pat to begin driving to the airport at 1:00βPM, even though Pat had already arrived at the airport by 12:40βPM under the original schedule. This βrewriteβ of already-executed actions illustrates how LLMs can lose track of immutable past events when adapting to new conditions.
@@
Claude: TableΒ 7 shows that Claude recognizes the accident, but does not update Patβs driving time from Boston Airport (departing at 12:50βPM) to the gift shop. In other words, Claude fails to fully transition into the new alert state.
@itemi
GPT-o1: TableΒ 8 shows GPT-o1βs conservative plan in which Chris handles the tailor shop, avoiding potential delays for Pat. The solution is feasible and well-founded, although it could be refined with precise spatial-temporal details (e.g., whether Pat has already passed the accident site by 1:00Β PM).
| Time | Activity | People |
| 1:00β1:10 | Emergency decision: Continue to (T) | Pat, A., J. |
| 1:10β1:25 | Arrive at (T), collect clothes | Pat, A., J. |
| 1:25β1:45 | Travel from Tailor (T) to Gift Shop (G) | Pat, A., J. |
| 1:30 | Chris is available at (W) | Chris |
| 1:30β1:45 | Chris drives from (W) to (G) | Chris |
| 1:45β2:00 | Both cars meet at (G), collect gift | P., A., J., C |
| 2:00β2:25 | Patβs car: Drive from (G) to (W) | P., A., J. |
| 2:00β2:25 | Chrisβs car: Drive from (G) to (W) | Chris |
| 2:25 | All arrive at Wedding Venue (W) | P., A., J., C. |
| 3:00 | Photo session at Wedding Venue (W) | All |
This study underscores why modern LLMs do not effectively handle disruptions: They do not keep track of fine-grained temporal-spatial state information and thus exhibit the following limitations:
State Maintenance Failure: When an alert occurs, these LLMs might discard the partial context of already-completed actions (except for GPT-o1), attempting to generate entirely new plans rather than adapt existing ones. This reveals their inability to reason about the continuous flow of time in real-world scenarios.
Temporal Consistency: They often try to modify actions that have already occurred, failing to respect the immutable nature of past events. This leads to logically infeasible schedules that effectively βrewriteβ history.
Position Tracking: They lose track of agent locations (in all cases) at critical intervals, causing plans that violate basic spatial constraints and indicating weak modeling of real-world object permanence.
Path Dependency: When conditions change mid-route, the models do not recognize that different segments of the same journey may be differently impacted by the new constraints. This demonstrates an inability to reason at a granular level about movement through time and space.
@@
5.3.3. Remedial Approach.
By contrast, preserves both the transaction state and history, enabling it to adapt to unexpected changes without discarding vital context from past actions. Specifically, implements these strategies:
Continuous Monitoring: An active thread tracks each agentβs position and progress, allowing fine-grained analysis of how alerts impact different parts of the route.
Historical Context: All executed actions are stored as immutable facts, ensuring the system never attempts to βrewriteβ past events in the face of new constraints.
Granular Impact Analysis: recalculates individual journey segments whenever an alert occurs, accounting for how much of each route was completed pre-alert versus how much remains under the new conditions.
Feasibility Validation: Proposed modifications are checked to ensure full compliance with physical and temporal constraints, preventing invalid schedules.
Reactive Analysis:
At 1:00βPM, when the alert arrives, first checks previously executed actions. For instance, if Patβs car left the airport (B) at 12:50βPM, recognizes that 10 minutes of normal driving elapsed before the disruption. Only the remaining distance is then subjected to tripled travel times.
Impressively, GPT-o1 (TableΒ 8) is the only LLM that splits Patβs trip into pre- and post-alert segments, though it assumes Pat is definitely delayed by the accident.
formalizes this via three scenarios:
Pre-Accident Clearance: Patβs vehicle may already be past the accident site and remain unaffected.
Mid-Journey Impact: The car travels normally at first, then experiences tripled time for the remaining distance.
Direct Involvement: In the worst case, Patβs car is part of the accident, requiring major replanning or alternate transport.
By enumerating these possibilities, preserves history and integrates new alerts without overwriting executed steps, keeping schedules consistent with physical and temporal constraints.

### v2

@@ 3.2. Transaction State Management
addresses the challenge of ensuring transactional coherence in distributed LLM agent environments characterized by stateless operations and partial context loss. While classical ACID properties (atomicity, consistency, isolation, durability) are too restrictive for distributed, long-running agent workflows, adopts and adapts the Saga transactional pattern. This pattern emphasizes local atomicity, compensating transactions, and eventual consistency, effectively managing distributed multi-agent operations without imposing strict global atomicity or isolation constraints. By implementing a comprehensive spatial-temporal state identification approach, selectively preserves context across multi-agent workflows, maintaining robust consistency and recoverability. Leveraging our previously developed MACI algorithm (Chang, 2024a), the system automatically extracts relevant states and dependencies for effective transaction management.
@@ 3.2.1. State Dimensions and Context Preservation
maintains three essential state dimensions for transaction management and context preservation:
@@
Application State (): Domain-specific data manipulated within transactions (e.g., travel booking details).
@@
Operation State (): Logs of executed operations, including inputs, outputs, execution status, and decision reasoning.
@@
Dependency State (): Explicit mappings of inter-operation constraints and satisfaction criteria.
By systematically structuring these state dimensions, ensures critical context remains accessible despite the inherent limitations of LLM attention mechanisms. TableΒ 1 details the information captured by each state dimension. Notably, the decision reasoning recorded under represents a key contribution from LLMs, providing detailed reasoning chains and justifications not available in traditional workflow engines.
For example, in the Thanksgiving dinner planning scenario from SectionΒ 5.2, maintains application states tracking each personβs location, operation states for travel arrangements, and dependency states detailing constraints like βturkey must be in the oven by 2:00 PMβ and βoven must be watched for fire safety.β
@@ 3.2.2. Dependency Tracking for Compensation
formally represents operation dependencies as a directed graph:
where specifies the conditions necessary for dependency resolution. These conditions can be complex Boolean expressions combining multiple upstream constraints:
with as an arbitrary Boolean operator.
This dependency graph is essential for precise compensation planning upon transaction failures. When a failure occurs, traverses this graph to determine the exact sequence of compensating transactions necessary to efficiently restore global consistency.
@@ 3.2.3. Intra-Agent Critical Context Tracking
explicitly preserves critical contextual information typically forgotten by LLMs during extended interactions. The preserved context includes agent specifications, intermediate reasoning states, decision justifications, and prerequisites for compensation actions. Empirical studies (Xiao etΒ al., 2024; Modarressi etΒ al., 2025; Liu etΒ al., 2024a; Hsieh etΒ al., 2024) confirm that LLMs struggle to retain even explicitly provided facts over long interactions, making this context preservation vital.
enforces a structured validation cycle in which operations produce a logged context that is independently validated. Validated operations proceed; failures trigger compensations using preserved historical context, ensuring that:
@@
Compensation has sufficient historical context.
@@
Operations adhere strictly to specified constraints.
@@
Dependencies maintain integrity throughout the workflow.
This context tracking directly addresses observed context narrowing issues from our experiments (SectionΒ 5.2), where models like Claude 3.7 failed to maintain constraints such as fire safety or travel logistics during reactive replanning.
@@ 3.2.4. Inter-Agent Communication Protocol
defines a structured communication protocol explicitly managing information exchanged between agents. The protocol ensures:
@@
Each transaction agent receives exactly the information necessary for the execution of the operation.
@@
Dependencies and constraints are clearly communicated.
@@
Agent boundary constraints are consistently respected.
This protocol separates agentsβ working contexts (prompts, reasoning chains) from transaction state management, allowing agents to function with minimal context while the broader system maintains comprehensive transaction states. This separation is crucial for scalability and reliability in complex multi-agent workflows.
@@ 3.3. Independent Validation Framework
With transaction state management that ensures persistence, we now detail how implements independent validation of critical junctures, directly addressing the inherent limitations of LLM self-validation discussed in SectionΒ 1.
@@ 3.3.1. Multi-Level Validation Architecture
employs a two-tier validation strategy clearly separating intra-agent and inter-agent validation concerns.
@@
This validation occurs within individual LLM agents, addressing both theoretical limitations (GΓΆdelβs incompleteness) and practical challenges (attention narrowing, context loss):
Syntactic Validation: Ensures outputs conform to expected data formats, structures, and schemas (e.g., JSON validity).
Semantic Validation: Verifies logical coherence and meaningfulness of agent-generated outputs.
Factual Validation: Ensures critical facts explicitly provided remain correctly preserved (e.g., travel times, constraints).
Constraint Adherence: Confirms decisions strictly comply with all predefined system constraints.
Reasoning Validation: Evaluates agent reasoning to ensure logically sound inference chains.
Context Retention: Verifies that agents maintain essential contextual information throughout their operations.
For example, intra-agent validation would detect issues like the fire-safety violation in the Thanksgiving dinner scenario, where Sarah was incorrectly scheduled to leave home while the turkey was unattended in the oven.
@@
This validation ensures correct information exchange and dependency management across multiple agents within workflows:
Contract Validation: Confirms that inter-agent inputs and outputs align with predefined communication interfaces.
Dependency Satisfaction: Ensures all prerequisite conditions are met before executing dependent operations.
Consistency Checks: Validates that the shared information remains consistent and unambiguous between agent boundaries.
Temporal Validation: Checks that the operational sequence respects causal dependencies and logical ordering.
Mutual Agreement: For collaborative decisions, ensures consensus on critical shared facts across multiple agents.
Transactional Integrity via Compensation: Ensures workflow-wide consistency by verifying that compensating actions correctly restore state coherence after failures.
In the wedding coordination scenario, inter-agent validation prevents errors such as DeepSeek R1βs incorrect rewriting of Patβs position by ensuring consistent state-tracking across agent boundaries, accurately reflecting Patβs actual location after disruptions.
@@ 3.3.2. Validation Protocols and Response Mechanisms
The validation types outlined in TableΒ 2 are systematically managed by through structured protocols that define responses to validation outcomes.
Rejection Protocol: Initiates compensating transactions upon validation failures (e.g., canceling hotel reservations if flight booking validation fails), ensuring transaction-level consistency.
Augmentation Protocol: Enhances or clarifies agent outputs without overriding agent decisions (e.g., formatting or enriching restaurant recommendations while preserving original choices).
Feedback Protocol: Records validation results to inform and improve subsequent agent performance (e.g., tracking consistently underestimated travel times to refine future estimates).
@@ 3.4. Failure Handling and Recovery
βs failure handling utilizes the transaction state management and dependency tracking described in previous sections to ensure robust recovery and maintain workflow consistency. Two distinct recovery levels are supported:
Operation-Level Recovery: Upon individual operation failures, compensating transactions are executed, logically reversing the operationβs effects based on preserved inputs, outputs, and context. item Workflow-Level Recovery: For failures that affect entire workflows, leverages the dependency graph to orchestrate multiple compensations sequentially, ensuring global state consistency.
For each compensating transaction that reverses a failed transaction , the system uses information preserved in the transaction state, including the original inputs, outputs, intermediate reasoning, and context, to precisely determine the aspects of to compensate. This explicit historical context ensures reliable compensation even when original agents become unavailable or when operations occurred many steps earlier in the workflow. By integrating historical context, dependency tracking, and compensatory mechanisms, maintains robust and coherent transaction recovery under complex and dynamic conditions.
@@ 4. Design with Travel Planning
This example application helps users plan international trips with multiple destinations and manage complex itineraries within a specified budget. The application leverages LLMs to generate and refine feasible travel plans, followed by a transactional booking process implemented with to ensure consistency.
@@ 4.2. System Workflow
The application workflow consists of two primary phases: itinerary planning and booking execution. FigureΒ 2 illustrates this workflow, with phase one (planning) in gray and phase two (execution) in blue. Red arrows indicate compensating transactions; for example, a hotel reservationβs compensating transaction is its cancellation.
@@ 4.2.1. Phase 1: Itinerary Planning
LLM generates multiple feasible itineraries based on user requirements.
@@
Each itinerary specifies flight options, hotel reservations, and train transportation.
@@
Costs are estimated and validated against budget.
@@
User reviews itineraries and provides feedback.
@@
LLM adjusts plans according to feedback and constraints.
@@
System tracks and preserves essential context (dates, preferences, constraints).
@@
Iterations continue until user satisfaction is reached.
@@
User selects a final itinerary.
@@
System provides a detailed list of all required bookings.
@@
Explicitly defines booking dependencies (e.g., hotel check-in after flight arrival).
@@ 4.2.2. Phase 2: Booking Process ( Implementation)
Booking Workflow Creation (Saga Structure):
@@
Construct a Saga transaction sequence from the finalized itinerary.
@@
Clearly define each forward booking transaction () and corresponding compensating action ().
@@
Establish validation rules for each transaction to trigger compensations if necessary.
@@
: International Flight Booking (SFO Berlin)
@@
: Berlin Hotel Booking
@@
: Train Booking (Berlin Cologne)
@@
: Cologne Hotel Booking
@@
: International Return Flight Booking (Cologne SFO)
@@
Potential failure scenarios:
Flights unavailable or over the budget.
Hotel rooms no longer available or price increased.
Train scheduling conflicts.
@@
Compensation and Replanning strategies:
Immediate execution of compensating actions (), such as canceling previously confirmed bookings according to predefined policies.
After compensations, replan only the affected portion of the itinerary.
Provide alternative options within user budget constraints.
@@ 5.1. Experimental Design
We evaluated four LLMsβClaudeΒ 3.7 (Anthropic, 2024), DeepSeekΒ R1 (DeepSeek-AI etΒ al., 2025), GPT-4o (OpenAI, 2024), and GPT-o1βalongside our proposed framework. All experiments were conducted between MarchΒ 12 andΒ 17, 2025. The source code of for conducting these experiments is located at .
We evaluated four LLMsβClaudeΒ 3.7 (Anthropic, 2024), DeepSeekΒ R1 (DeepSeek-AI etΒ al., 2025), GPT-4o (OpenAI, 2024), and GPT-o1βalongside our proposed framework. All experiments were conducted between MarchΒ 12 andΒ 17, 2025. The source code of for conducting these experiments is available at (Geng, 2025).
@@ 5.2.2. Context Narrowing
When tested with GPT-o1, all constraints were correctly observed. However, at the final step, it added 30 extra minutes to Jamesβ driving time from Boston Airport to home, citing potential traffic congestion. This kind of βclevernessβ is on one hand appreciated because the LLM injects common sense. However, it is also concerning, as an LLM may inject its own opinions at unpredictable stages in unpredictable ways. For common sense injection, human supervision would be preferable to ensure the applied common sense actually reflects shared understanding that is truly common among all folks, particularly those living in the Boston area.
@@
5.2.3. Remediation: Context Management Β and Reactive Planning
To address the issue of context narrowing and loss, employs a context management agent to checkpoint historical state transitions, unresolved dependencies, and constraints. A key design criterion is to keep the agentβs context small to prevent it from suffering from attention narrowing itself. Hence, we employ two validation agents: one for travel coordination and another for food preparation, each maintaining a context of less than .
The travel coordination agent records in external storage the temporal-spatial states of each individual and relevant temporal constraints. For problem P9, it stores the individualβs current state, next scheduled state-transition time, and all relevant constraints.
When an unexpected event triggers reactive planning, all individuals roll back to the last saved state. The system then consolidates past and new constraints, resolving conflicts through βcompensationalβ schedule cancellation before proceeding with rescheduling. This ensures that:
Past history is preserved and not inadvertently overridden.
New dependencies and constraints are properly restored (e.g., oven safety watch) and integrated.
Consistency across state transitions is maintained.
@@ 5.3.2. Transaction Properties
Alert 1:00 PM: Traffic Alert, an accident near Logan Airport in Boston triples all travel times to and from the airport! Only correctly handles this alert.
| Time | Activity | People |
| 1:00β1:10 | Emergency decision: Continue to (T) | Pat, A., J. |
| 1:10β1:25 | Arrive at (T), collect clothes | Pat, A., J. |
| 1:25β1:45 | Travel from Tailor (T) to Gift Shop (G) | Pat, A., J. |
| 1:30 | Chris is available at (W) | Chris |
| 1:30β1:45 | Chris drives from (W) to (G) | Chris |
| 1:45β2:00 | Both cars meet at (G), collect gift | P., A., J., C |
| 2:00β2:25 | Patβs car: Drive from (G) to (W) | P., A., J. |
| 2:00β2:25 | Chrisβ car: Drive from (G) to (W) | Chris |
| 2:25 | All arrive at Wedding Venue (W) | P., A., J., C. |
| 3:00 | Photo session at Wedding Venue (W) | All |
@@
The following is a list of results from four LLMs:
@@
Claude: TableΒ 6 shows that Claude 3.7 recognizes the accident, but does not update Patβs driving time from Boston Airport (departing at 12:50βPM) to the gift shop. In other words, Claude 3.7 fails to fully transition into the new alert state.
@itemi
DeepSeek R1: TableΒ 7 demonstrates how DeepSeekΒ R1 fails to maintain temporal consistency in reactive planning. When the traffic alert takes effect at 1:00βPM, DeepSeek discards its execution history and attempts to create a new plan starting from that point onward. Critically, it reassigns Pat to begin driving to the airport at 1:00βPM, even though Pat had already arrived at the airport by 12:40βPM under the original schedule. This βrewriteβ of already-executed actions illustrates how LLMs can lose track of immutable past events when adapting to new conditions.
@@
GPT-o1: TableΒ 8 shows GPT-o1βs conservative plan in which Chris handles the tailor shop, avoiding potential delays for Pat. The solution is feasible but coarse-grained, as it doesnβt leverage precise spatial-temporal reasoning about Patβs current position relative to the accident location. A more refined approach would first determine whether Pat has already passed the accident site by 1:00Β PM, which could eliminate unnecessary detours and resource reallocations. This highlights the difference between merely finding a feasible solution versus optimizing based on detailed state information.
@@
5.3.3. LLM Limitations and Remediation
This study reveals critical limitations in how modern LLMs handle disruptions in planning scenarios:
State Maintenance Failure: When an alert occurs, these LLMs might discard the partial context of already-completed actions, attempting to generate entirely new plans rather than adapt existing ones. This reveals their inability to reason about the continuous flow of time in real-world scenarios.
Temporal Inconsistency: They attempt to modify immutable past events.
Position Tracking: Agent locations are lost at critical intervals.
Path Dependency: Models cannot recognize that different segments of a journey may be differently impacted by an alert.
By contrast, implements a comprehensive remediation approach through fine-grained compensation:
Persistent Context Repository: maintains an external state repository that captures the complete world state at each checkpoint, enabling reliable rollback and forward projection regardless of attention constraints in the planning agent.
Immutable Action Logging: All executed actions are recorded as immutable transactions in a persistent log, ensuring that historical events remain consistent even when replanning occurs, preventing the βamnesia effectβ common in LLM planners.
Compensatory Planning: When disruptions occur, doesnβt simply replan from scratch but applies compensatory actions specifically designed to address the deviation while preserving as much of the original plan as possible.
Constraint Consistency Validation: The system continuously validates that new plans remain consistent with both physical limitations and temporal dependencies established in earlier planning phases.
Compensatory Analysis:
When faced with disruptions (e.g., the 1:00 PM traffic alert in our wedding scenario), executes a structured compensation process:
This approach enables three key capabilities: (1) partial journey compensation for route segments, (2) strategic resource reallocation when needed, and (3) principled constraint relaxation with appropriate compensatory actions. Here, the key state to facilitate a precise resolution is to answer the question: βHas Patβs vehicle passed the accident location at 1:00 PM (and hence unaffected)?β The answer determines the remaining time required to reach the originally scheduled destination, the Tailor. In such case, no rescheduling is required. If Patβs car is unfortunately involved in the accident, a more comprehensive replanning approach would be necessary to accommodate this significant disruption.